# Machine Vision-Based Assessment of Fall Color Changes and its Relationship with Leaf Nitrogen Concentration

### v2

@@ 4 Conclusion and Future Work
Leaf Nitrogen concentration is one of the critical factors that determines the yield and quality of fruits and the overall health of the apple trees. Growers often rely on different visual cues including the canopy color for assessing plant health and nitrogen needs. This study presented a machine vision-based approach for quantitatively assessing tree canopy color and utilizing the color during transitional period in the fall (when the leaves change color from green to yellow during senescence) to assess N status in trees. Specifically, this study segmented the test tree canopies in an outdoor environment with natural background using 3D point clouds, identified yellow and green regions in the canopy using a gradient boost model, and used a metric called yellowness index to quantify the canopy color. It was found that leaf nitrogen concentration in apple trees was significantly correlated to the yellowness index of the tree canopies at the optimal color change stage in the fall. The trees notably began transitioning in color around the 29th week following full bloom in both years the test was conducted. The trees with lower and higher nitrogen showed the significant differences between the yellowness indices during week 4 (Around November 4) in 2021 and week 2 (October 27) and week 4 (November 10) in 2023. The difference in these dates signifies the importance of factors such as temperature, humidity, and growing degree days during different years in predicting optimal window of estimate yellowness index for tree nitrogen levels assessment. Overall, the findings from this study could be summarized as:
@@
    A machine vision-based system could be used to segment the yellow and green foliage in apple trees in outdoor orchard environments. A metric yellowness index was defined to quantify the status of tree foliage color, which was estimated with a $R^{2}$ value of 0.72 using a gradient boost classifier model .
@@
    The yellowness index of the trees was found to be correlated with the leaf nitrogen levels in the trees. Trees with different nitrogen levels showed different yellowing patterns. The difference between low and high nitrogen trees were more significant ($p<0.05$) in week 4 in 2021(31 weeks after full bloom) and 2023 (29 weeks after full bloom).
This study presented a novel way of assessing the color of apple tree canopies and explored its relationship with leaf Nitrogen concentration. This method could be a good alternative to traditional methods like chemical methods, chlorophyll meter, and spectral analysis of N assessment in apple trees and can give instantaneous results that could be expanded to individual tree level assessment using a ground vehicle or robot. The study also provided critical insights into the fall color changes in apple trees and their relationship with leaf nitrogen concentrations.
@@
For practical implementation, machinery capable of delivering nutrients at a per-tree level becomes essential. Several companies and academic institutions are developing autonomous systems for precise orchard operations (Brown et al., 2024). Integrating the findings from this study into such machinery would provide a comprehensive solution for growers to apply right amount of N to individual trees thus substantially enhancing tree health, and fruit yield and quality, and ultimately increase returns while insuring long term sustainability of apple production.

### v3

@@ 4 Conclusion and Future Work
Leaf Nitrogen concentration is one of the critical factors that determines the yield and quality of fruits and the overall health of apple trees. Growers often rely on different visual cues including the canopy color for assessing plant health and nitrogen needs. This study presented a machine vision-based approach for quantitatively assessing tree canopy color and utilizing the color during transitional period in the fall (when the leaves change color from green to yellow during senescence) to assess N status in trees. Specifically, this study segmented the test tree canopies in an outdoor environment with natural background using 3D point clouds, identified yellow and green regions in the canopy using a gradient boost model, and used a metric called yellowness index to quantify the canopy color. It was found that the trees notably began transitioning in color around the $29^{t}h$ week following full bloom in both years the test was conducted. The trees with lower and higher nitrogen showed significant differences between the yellowness indices during week 4 (Around November 4) in 2021 and week 2 (October 27) and week 4 (November 10) in 2023. The difference in these dates signifies the importance of factors such as temperature, humidity, and growing degree days during different years in predicting optimal window of estimate yellowness index for tree nitrogen levels assessment. Overall, the findings from this study can be summarized as:
@@
    A machine vision-based system can be used to segment yellow and green foliage in apple trees in outdoor orchard environments. A metric yellowness index was defined to quantify the status of tree foliage color, which was estimated with a $R^{2}$ value of 0.72 using a gradient boost classifier model.
@@
    The yellowness index of the trees was found to be correlated with the leaf nitrogen levels in the trees during some periods of the study. Trees with different nitrogen levels showed different yellowing patterns. The difference between low and high nitrogen trees were more significant ($p<0.05$) in week 4 in 2021(31 weeks after full bloom) and 2023 (29 weeks after full bloom).
This study presented a new approach for assessing the color of apple tree canopies and explored its relationship with leaf Nitrogen concentration. This method could be a good alternative to traditional methods like chemical methods, chlorophyll meter, and spectral analysis of N assessment in apple trees and can give instantaneous results that could be expanded to individual tree level assessment using a ground vehicle or robot. The study also provided critical insights into the fall color changes in apple trees and their relationship with leaf nitrogen concentrations.